What is the longest wavelength of a photon that can cause the dissociation of the O-O bond in ozone ?

Type your answer here: nm
Answer #1

Answer-  Bond is O-O

Longest wavelength required to break this bond = ?

Energy required to break 1 mol of O-O bond is 495 kJ / mol (known value determined experimentally)

so, energy required to break 1 molecule of 0-O bond is 495 kJ /Na(Avogadros no.) i.e

495/ 6.022×10+23 = 82.2 * 10-23 kJ = 82.2 * 10-20 J ( 1 kJ = 1000J)

We know that,

E = hc/\lambda

where, E = Energy

h = plank's constant = 6.63×10−34 m2 kg / s

\lambda = wavelength

c = speed of light = 3 * 108 m/s

Put all the values,

E = hc/\lambda

82.2 * 10-20 = (6.63×10−34 * 3 * 108 ) / \lambda

\lambda = (6.63×10−34 * 3 * 108 ) / 82.2 * 10-20

\lambda = 0.241971 * 10-6 m = 241.971 * 10-9 m

1 m = 10-9 nm

So, 241.971 * 10-9m will be equal to 241.971 * 10-9 * 10-9 nm = 241.97 nm

longest wavelength required to for dissociation is 241.97 nm.